OpenStudy (anonymous):

theses are fractions.... 5/6x-1/5x+4=7/10x

OpenStudy (jack1):

yes... they are ;D

OpenStudy (anonymous):

lol haha last problem that had fractions someone didnt realize that was what they was

OpenStudy (jack1):

\[ \frac{ 5 }{ 6x } - \frac{ 1 }{ 5x } + 4 = \frac{ 7 }{ 10x }\] is the above right?

OpenStudy (anonymous):

yeah how did u get the line under it lol

OpenStudy (jack1):

if u tap the equation button below the chatbox here, it opens up a math screen thingy

OpenStudy (jack1):

anyways: easiest solution: flip it

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 }x\]

OpenStudy (anonymous):

aww i see lol thank you very helpful

OpenStudy (jack1):

\[\frac{ 5 }{ 6x }−\frac{ 1 }{ 5x }+4=\frac{ 7 }{ 10x } \rightarrow \frac{ 6x }{ 5 } - \frac{ 5x }{ 1 } +\frac{ 1 }{ 4 } = \frac{ 10x }{ 7 }\]

OpenStudy (jack1):

nah, no worries ;D

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

one from last night \[\frac{ 4 }{ 5}x-\frac{ 1 }{ 4 }x+14=\frac{ 9 }{ 10 }x\]

OpenStudy (anonymous):

and for last night i got 40?

OpenStudy (jack1):

ahh, sorry, my bad

OpenStudy (jack1):

cool, so all u need is common denominator and you can add the tops together

OpenStudy (jack1):

(tops = numerators)

OpenStudy (jack1):

so common denominator of 1/5 , 1/6 and 1/10 = 1/30 so 5/6 becomes 25 / 30 1/5 becomes 6/30 7/10 becomes 21/30 new equation = \[\frac{ 25 }{ 30 }x -\frac{ 6 }{ 30 }x + 4 = \frac{ 21 }{ 30 }x\]

OpenStudy (jack1):

now just add / subtract like terms as needed

OpenStudy (jack1):

@KupKaKe89 ? this all making sense?

OpenStudy (anonymous):

mmm no lol i dont no anything about fractions

OpenStudy (anonymous):

i got 60

OpenStudy (jack1):

that's cool, fractions are just a way of thinking about decimals so 0.75 = three divided by four also equals 12 divided by 16 also equals 120 divided by 160 and can be represented by the fractions : 3/4 or 6/8 or 9/12 ... etc it all still refers to the one number (0.75)

OpenStudy (jack1):

so with fractions, what you do to the top number, you must do to the bottom number (otherwise its not referring to the same number) so if we want to add 1/6, 1/7 and 3/4 together, we need a common base number (denominator) current denominators are 6, 7 and 4 so want's one number all are factors of? 6 x 7 x 4 = 168 so 168 is the common denominator 1/6 = ? / 168 so 6 x what = 168 (7x4) so 6 x 7 x 4 = 168 therefore what you do to the bottom line u must do to the top line so 1 x 7 x 4 = 28 so new fraction = 28 / 168 etc...

OpenStudy (jack1):

and yeah, x = 60